OpenStudy (anonymous):

The coordinates below represent two linear equations. How many solutions does this system of equations have? Line 1 x y –1 6 0 4 Line 2 x y 0 7 1 5 A.0 B.exactly 1 C.exactly 2 D.infinitely many please help. will fan and medal :)

OpenStudy (texaschic101):

line 1...slope = (y2 - y1) / (x2 - x1) slope = (4 - 6) / (0 - (-1) slope = -2/1 slope = -2 y = mx + b 6 = -2(-1) + b 6 = 2 + b 4 = b equation is : y = -2x + 4 ------------------ line 2 slope = (5 - 7) / (1 - 0) slope = -2/1 slope = -2 y = mx + b 7 = -2(0) + b 7 = b equation is : y = -2x + 7 ------------------- if they have same slope but different y intercepts, then the lines are parallel and do not cross. Therefore, they have 0 solutions

OpenStudy (anonymous):

ohh. i see now. thanks :)

OpenStudy (texaschic101):

YW :)